# Cooperative Payload Estimation by a Team of Mocobots

### v1

@@ IV-B Data Processing
Figure 4: Frame $\{i\}$ of the 6-dof manipulator of Omnid $i$ is positioned at the center of the gimbal wrist.
Let $\theta=(\theta_{1},\theta_{2},\theta_{3})$ represent the actuated proximal joint angles of the Delta mechanism and $\phi=(\phi_{x},\phi_{y},\phi_{z})$ represent the angles of the gimbal wrist, all measured by encoders. The transformation matrix $T_{i_{0}i}$ can be derived through forward kinematics as
$T_{i_{0}i}(\theta,\phi)=\begin{bmatrix}g(\phi)&h(\theta)\\ 0&1\end{bmatrix}.$ (34)
@@
The wrench at $\{i\}$ is ${\mathcal{F}}_{i}=(0,f_{i})$, where $f_{i}=(\partial h/\partial\theta)^{-\intercal}\tau_{i}$ and $\tau_{i}$ is the Delta’s SEA joint torque vector.
The derivative $\dot{T}_{i_{0}i}$ and the twist derivative $\dot{{\mathcal{V}}}_{i}$ are calculated using central differencing post-processing of the $100$ Hz estimates of $T_{i_{0}i}$ and ${\mathcal{V}}_{i}$, respectively. To smooth the data and reduce the effect of amplification of encoder quantization in this numerical differencing, we apply a fourth-order Butterworth low-pass filter to the encoder data prior to calculating the $T_{i_{0}i}$ estimates. The same filter is applied to smooth the SEA torque data.
@@ IV-C Experimental Results
IV-C1 Transformation Matrix
The payload was moved randomly and manually in all 6-dof by a human operator using the float mode of the Omnid [1], a mode compensating for gravitational forces using active force control. The encoder data from the Omnid manipulators were collected to calculate ${\mathcal{V}}_{i*},i=1,\ldots 3$. The transformations for the individual robot pairs $\hat{T}_{12}$, $\hat{T}_{23}$, and $\hat{T}_{31}$ were first calculated using the least-squares methods of Section III-B, and these estimates were then refined using the nonlinear optimization incorporating loop-closure constraints. (In our experiments, this second optimization step did not change the estimates significantly.) Five independent trials were conducted, each consisting of ten seconds of data (1000 samples) for the estimation process.
@@ IV-C1 Transformation Matrix
Parameters Absolute Error
@@
Parameters Absolute Error Percentage Error
@@ IV-C2 Mass and Center of Mass
To estimate the mass and center of mass of the payload, the Omnids held the payload stationary at four different orientations. The results of the least-squares estimates from Section III-C are summarized in Table IV.
Parameters Absolute Error Percentage Error
mass (kg) $0.099$ $0.867$%
$p_{1c}$ (m) $0.026$ $4.403$%
@@ IV-C3 Inertia Matrix
To estimate the inertia matrix, the Omnids performed four one-minute manipulation trials, with the first and last four seconds of data discarded (5200 samples for each trial). In each manipulation, two Omnid wrists moved in periodic trajectories on the surface of the sphere for which the origins of the three grasp frames lie on a common meridian. The third Omnid had its actuators locked and passively accommodated the motion of the other two Omnids via its gimbal wrist and elastic Delta joints.
Table V summarizes the results from the four experimental trials, with the estimated inertia matrix transformed into the principal axes of the payload. Similar to the labeling of axes in the ground truth inertia matrix, the labels ^$\hat{{\mathcal{I}}}_{xx}$, ^$\hat{{\mathcal{I}}}_{yy}$, and ^$\hat{{\mathcal{I}}}_{zz}$ are given to principal inertia estimates in ascending order.
Parameters Absolute Error Percentage Error
@@
[…]
TABLE V: Experimental results of inertia matrix estimation.
V Discussion
The experimental results demonstrate the validity of the approach, producing relative grasp rotation matrix errors typically less than $3^{\circ}$ and position vector magnitude errors of less than $4\%$, a mass estimate error of less than $1\%$, and a CoM position error of less than $5\%$. Inertia matrix estimation, which relies on acceleration estimates from twice-differenced encoder data, exhibits a bit more error, but it is still less than $10\%$ for the principal moments of inertia.
@@ V Discussion
First, grasp frame velocities and accelerations are obtained by differencing encoder data, including encoders at the proximal joints of the Delta mechanism, distant from the wrist. The intervening mechanism and the differencing process introduce noise to the estimation. To address this, the Omnid sensor suite could be supplemented by IMUs at the gimbals. Independent of the sensing modality, however, the inherent compliance of the manipulators, due to their SEAs, can introduce vibrations at the wrists, which may be exacerbated by a poor choice of wrist trajectories to estimate the payload inertial properties. Because the Omnids are designed primarily for force-controlled human interaction, they are incapable of stiff, highly precise motion control, which is assumed to be supplied by the human during typical human-robot collaboration.
Second, manipulation forces at the wrist are estimated based on the Delta’s proximal joint SEAs. This assumes zero friction at joints and bearings, but eliminates the need for costly and potentially fragile end-effector force-torque sensors.
@@ VI Conclusion
We present a methodology for cooperative payload estimation using a group of mocobots, focusing on the estimation of the mass, center of mass, and inertia matrix of a rigid body payload, as well as the transformations between the grasp frames of each robot. The proposed method avoids the need for expensive sensors and intercommunication between robots by leveraging the twists, twist derivatives, and wrenches measured at each robot’s grasp frame. Despite challenges of sensor noise and unmodeled dynamics, the proposed approach demonstrates promising results for payload estimation using Omnid mocobots. Future work will focus on extending the methodology to identify the kinematic and inertial properties of articulated payloads, and eventually continuously deformable payloads.

### v3

@@ IV-B Data Processing
Figure 6: Frame $\{i\}$ of the 6-DoF manipulator of Omnid $i$ is positioned at the center of the gimbal wrist. The $\hat{z}$-axis encoder is installed at the base of the gimbal wrist, which is not visible in the figure. The left image shows the manipulator in its home position.
As shown in Figure 6, $\phi=(\phi_{1},\phi_{2},\phi_{3})$ represent the actuated proximal joint angles of the Delta mechanism and $\theta=(\theta_{x},\theta_{y},\theta_{z})$ represent the angles of the gimbal wrist, all measured by encoders. The transformation matrix $T_{i_{0}i}$ can be derived through forward kinematics as
$T_{i_{0}i}(\theta,\phi)=\begin{bmatrix}g(\theta)&h(\phi)\\ 0&1\end{bmatrix}.$ (35)
@@
The wrench at $\{i\}$ is ${\mathcal{F}}_{i}=(0,f_{i})$, where $f_{i}=(\partial h/\partial\phi)^{-\intercal}\tau_{i}$ and $\tau_{i}$ is the Delta’s SEA joint torque vector.
The derivative $\dot{T}_{i_{0}i}$ and the twist derivative $\dot{{\mathcal{V}}}_{i}$ are calculated using central differencing post-processing of the 100100100 Hz estimates of $T_{i_{0}i}$ and ${\mathcal{V}}_{i}$, respectively. To smooth the data and reduce the effect of amplification of encoder quantization in this numerical differencing, we apply a third-order Butterworth low-pass filter to the encoder data prior to calculating the $T_{i_{0}i}$ estimates. The same filter is applied to smooth the SEA torque data. To avoid filter edge effects and other transients, the first and last two seconds of each trial’s data are discarded.
@@ IV-C Experimental Results
IV-C1 Grasp Kinematics
To generate sufficient twist information for accurate estimation of relative grasp frame locations, the payload was randomly moved in all six DoF by assigning independent random trajectories to each Omnid manipulator. Each random trajectory consisted of a series of nominally straight-line paths for the gimbal wrist between via points, with the randomized transit times and dwell times at the vias chosen to be between 0.50.50.5 and 0.80.80.8 s. Because these independent random trajectories are not kinematically compatible with rigid grasps on the payload, we took advantage of mechanical compliance and used low motion control gains at each Omnid’s force-controlled manipulator to allow trajectory tracking error to keep internal forces on the payload low.
Encoder data from the Omnid manipulators were collected to calculate ${\mathcal{V}}_{i*},i=1,2,3$. The transformations for the individual robot pairs $\hat{T}_{12}$, $\hat{T}_{23}$, and $\hat{T}_{31}$ were first calculated using the least-squares methods of Section III-B, and these estimates were then refined using nonlinear optimization incorporating loop-closure constraints. In our experiments, this second optimization step did not change the estimates significantly. Six independent trials were conducted for each payload, and the random trajectories for each trial consisted of 808080 random via points.
@@ IV-C1 Grasp Kinematics
Config.
Parameters
Absolute Error (deg)
@@
Config.
Parameters
Absolute Error (m) Percentage Error
@@ IV-C2 Mass and Center of Mass
To estimate the mass and center of mass of the payload, the Omnids were commanded to hold the payload stationary at six distinct orientations. Data collection was performed during periods when the measured forces along all axes remained constant (with a tolerance of 0.010.010.01 N) for a continuous duration of six seconds. The results of the least-squares estimates from Section III-C are summarized in Table IV.
Config. Parameters Absolute Error Percentage Error
(a) ${\mathfrak{m}}$ (kg) 0.11 0.9%
$p_{1c}$ (m) 0.021 3.7%
(b) ${\mathfrak{m}}$ (kg) 0.07 0.6%
$p_{1c}$ (m) 0.018 3.3%
(c) ${\mathfrak{m}}$ (kg) 0.11 1.1%
$p_{1c}$ (m) 0.039 4.5%
(d) ${\mathfrak{m}}$ (kg) 0.18 1.8%
$p_{1c}$ (m) 0.023 3.8%
@@ IV-C3 Inertia Matrix
To estimate the inertia matrix, the Omnids were tasked with generating motion to sufficiently excite the payload’s rotational dynamics. Two types of trajectories were employed: the random trajectories used previously in the grasp kinematics experiments, and periodic trajectories constructed based on the estimated grasp transformations. For the periodic trajectories, all three Omnid wrists moved along periodic trajectories on the surface of the sphere for which the origins of the three grasp frames lie on a common meridian. In total, six experimental trials for each configuration were conducted: three trials using random trajectories, each consisting of 80 configurations per Omnid, and three trials using controlled periodic trajectories, each lasting one minute. No significant difference in the data quality was observed between the two trajectory types.
Table V summarizes the results from the six experimental trials. Similar to the labeling of axes in the ground truth inertia matrix, the labels ^$\hat{{\mathcal{I}}}_{xx}$, ^$\hat{{\mathcal{I}}}_{yy}$, and ^$\hat{{\mathcal{I}}}_{zz}$ are given to the axes with principal inertia estimates in ascending order.
Config.
Parameters
[…]
@@
(a) ${\mathcal{I}}_{xx}$ 0.083 0.036 3.6% 1.6%
${\mathcal{I}}_{yy}$ 0.130 0.022 4.1% 0.7%
${\mathcal{I}}_{zz}$ 0.102 0.166 1.8% 3.0%
(b) ${\mathcal{I}}_{xx}$ 0.067 0.065 3.1% 3.0%
${\mathcal{I}}_{yy}$ 0.086 0.102 2.5% 3.0%
${\mathcal{I}}_{zz}$ 0.246 0.056 4.4% 1.0%
(c) ${\mathcal{I}}_{xx}$ 0.005 0.073 0.3% 4.0%
${\mathcal{I}}_{yy}$ 0.039 0.120 1.6% 4.9%
${\mathcal{I}}_{zz}$ 0.158 0.091 3.8% 2.2%
(d) ${\mathcal{I}}_{xx}$ 0.048 0.058 2.8% 3.4%
${\mathcal{I}}_{yy}$ 0.016 0.060 0.8% 2.8%
${\mathcal{I}}_{zz}$ 0.005 0.165 0.1% 4.4%
TABLE V: Experimental results for the inertias of the principal axes of inertia for different payloads.
V Discussion
@@
First, grasp frame velocities and accelerations are obtained by differencing encoder data, including encoders at the proximal joints of the Delta mechanism, distant from the wrist. The intervening mechanism and the differencing process introduce noise to the estimation. To address this, the Omnid sensor suite could be supplemented by IMUs at the gimbals. Independent of the sensing modality, however, the inherent compliance of the manipulators, due to their SEAs, can introduce vibrations at the wrists, which may be exacerbated by a poor choice of wrist trajectories to estimate the payload inertial properties. In practice, our excitation trajectories resulted in relatively smooth motions of the payload, as shown in the video.
Second, manipulation forces at the wrist are estimated based on the Delta’s proximal joint SEAs. This measurement assumes zero friction at joints and bearings of the Delta mechanisms, but it eliminates the need for costly and potentially fragile end-effector force-torque sensors. Friction in the Omnid Delta joints is quite low, resulting in good wrist force sensing.
@@ VI Conclusion
We present a methodology for cooperative payload estimation using a group of mocobots, focusing on the estimation of the mass, center of mass, and inertia matrix of a rigid body payload, as well as the transformations between the grasp frames of each robot. Despite challenges of sensor noise and unmodeled dynamics, the proposed approach demonstrates promising results for payload estimation using Omnid mocobots. Future work will focus on extending the methodology to identify the kinematic and inertial properties of articulated payloads, and eventually continuously deformable payloads.